# Strong Data Processing Properties of Rényi-divergences via Pinsker-type Inequalities

### v2

@@ Abstract
We investigate strong data processing inequalities (SDPIs) of the Rényi-divergence between two discrete distributions when both distributions are passed through a fixed channel. We provide a condition on the channel for which the DPI holds with equality given two arbitrary distributions in the probability simplex. Motivated by this, we examine the contraction behavior for restricted sets of prior distributions via $f$-divergence inequalities: We provide an alternative proof of the optimal reverse Pinsker’s inequality for Rényi-divergences first shown by Binette [1]. We further present an improved Pinsker’s inequality for Rényi-divergence based on the joint range technique by Harremoës and Vajda [2]. The presented bound is tight whenever the value of the total variation distance is larger than $\nicefrac{{1}}{{\alpha}}$. By framing these inequalities in a cross-channel setting, we arrive at SDPIs that can be adapted to use-case specific restrictions of input distribution and channel. We apply these results to the Rényi local differential privacy amplification through post-processing by channels that satisfy no local differential privacy guarantee.
@@ I Introduction
[…]
$D_{\alpha}(P||Q)=\frac{1}{\alpha-1}\log\mathbb{E}_{Q}\biggl{[}\biggl{(}\frac{dP}{dQ}\biggr{)}^{\alpha}\biggr{]}$ (1)
@@
in an AWGN channel is derived in [10] via the Hockey-stick divergence $E_{\gamma}$. Recently, Jin et al. [11] show that the usual upper bound on contraction via total variation fails to hold for Rényi-divergences. They however prove that the universal $\chi^{2}$ lower bound contraction holds for general channels. Zamanlooy and Asoodeh [6] show that if a channel satisfies local differential privacy (LDP), SDPIs can be derived in terms of a channels LDP guarantee. These bounds are further applied to mixing times in [12]. Related, it is shown in [13] that the LDP guarantee of a channel is closely related to its $f$-divergence contraction properties.
Information-theoretic tools for assessing privacy in data processing systems have highlighted the crucial role of the Rényi-divergence in this context: On one hand, Rényi differential privacy (RDP) [14] has recently been adopted as a powerful relaxation of differential privacy. Since, it has been shown to be highly useful for, e.g, composition and privacy accounting [15, 16]. On the other hand, information-theoretic measures that aim to quantify privacy in terms of inference limits can often be shown to reduce to formulations involving Rényi-divergences. An early example of this is the usage of Rényi-divergence of order infinity in the framework of quantitative information flow [17, 18]. More recently, the highly related notion of maximal leakage [19] and its variants and generalizations [20, 21, 22, 23, 24] emerged, and solidify the important role of Rényi-divergence for adversarial approaches to privacy. Interestingly, the framework suggested in [25] unifies many of the above mentioned approaches to quantifying privacy, including maximal leakage, RDP and, in a local setting, Rényi local differential privacy (RLDP). On a related note, the recently proposed pointwise adaption of maximal leakage, pointwise maximal leakage (PML) [21] is shown in [26] to enable inferential guarantees that determine the privacy risk of outcomes by evaluating the Rényi-divergence of order $\infty$ between the prior and posterior distribution of the secret.
Rényi-divergence is itself not an $f$-divergence. However, it can be shown that the Rényi-divergence of any order $\alpha$ can be written as a monotone transform of an $f$-divergence parameterized by $\alpha$ [27]. Using this result, standard $f$-divergence inequalities can be applied to Rényi-divergence. Numerous such inequalities are derived in [27] and subsequent works. Perhaps closest to the approach used in this work, such inequalities are used to derive SDPIs in [28].
While the contraction of $f$-divergences–and hence of Rényi-divergences–of LDP channels is studied in works like [6, 29, 30, 31], there are fewer results that consider the contraction of channels that satisfy no LDP guarantee. For such channels, the aforementioned PML framework continues to allow for a (finite) privacy quantification. Since both frameworks rely on randomization for privacy protection, it is reasonable to assume that PML exhibits a similar relationship between contraction and privacy guarantee as observed for LDP.
Contributions: In this work, we derive SDPI-style bounds for Rényi-divergences aimed at information-theoretic privacy problems. We present a condition on the channel under which it does not contract the Rényi-divergence in general in Section III. In Section IV, we present tools for bounding Rényi-divergences in terms of total variation distance: We present an improved (and often tight) Pinsker’s inequality for Rényi-divergences given $\alpha>1$, as well as an alternative proof to the optimal reverse Pinsker’s inequality for Rényi-divergence first shown in [1]. We apply these tools to obtain results on the data processing properties of Rényi-divergence for discrete channels in Section V. The results show that the restriction of the input distributions can strongly improve the ability to obtain meaningful bounds on contraction. Finally, we illustrate the results by providing an example use-case that shows how these bounds enable an improved assessment of the RLDP privacy amplification through post-processing. The privacy properties of the non-RLDP channel in this bound are quantified by its PML. This dependence establishes a connection between the contraction of Rényi-divergence and the PML privacy parameter. Further, it shows how privacy assessment with LDP and PML can be used synergistically.
@@ II Preliminaries
We use uppercase letters to denote random variables, lowercase letters to denote their realizations and caligraphic letters to denote sets. Specifically, we consider the two random variables $X$ and $Y$, where $X$ represents input data into a transition kernel $P_{Y|X}$ (also referred to as channel or mechanism in the following) that induces a random variable $Y$ at its output. Let $P_{XY}$ denote the joint distribution of $X$ and $Y$. Then we use $P_{XY}=P_{Y|X}\times P_{X}$ to imply that $P_{XY}(x,y)=P_{Y|X=x}(y)P_{X}(x)$ and $P_{Y}=P_{Y|X}\circ P_{X}$ to denote the marginalization $P_{Y}(y)=\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X}(x)$. We write $\mathcal{P}_{\mathcal{X}}$ for the probability simplex on the set $\mathcal{X}$. For $N\in\mathbb{N}$, we define $[N]\coloneqq\{1,\dots,N\}$ as the set of positive integers up to $N$. Finally, $\log(\cdot)$ denotes the natural logarithm.
@@ II-A Rényi-divergence, $f_{\alpha}$-divergene and $E_{\gamma}$-divergence
We will use the Hockeystick-divergence $E_{\gamma}$ (from here on simply called $E_{\gamma}$-divergence) as a tool for obtaining bounds on the Rényi-divergence, as similarly done in [10, Section 3.2].
@@
can then be used to obtain any such $D_{f}$-divergence. For $\gamma=1$, the $E_{\gamma}$-divergence is equal to the total variation distance between the two distributions, i.e., $E_{1}(P||Q)=\text{TV}(P||Q)$.
@@ II-B Strong Data Processing Inequalities
All $f$-divergences (including the $f_{\alpha}$-divergence)–and hence also the Rényi-divergence–satisfy the data processing inequality (DPI). The DPI states that for any two probability measures $P_{X},Q_{X}$ on $\mathcal{X}$, and some transition kernel $P_{Y|X}$,
@@
We define the contraction coefficient $\eta_{\alpha}(P_{Y|X})$ of the $\alpha$-Rényi-divergence given a channel $P_{Y|X}$ as the maximum factor between the two sides of the above inequality.
@@
$=$ $\sup_{Q_{X}\in\mathcal{P}:D_{\alpha}(P_{X}||Q_{X})\neq 0}\frac{D_{\alpha}(P_{Y|X}\circ P_{X}||P_{Y|X}\circ Q_{X})}{D_{\alpha}(P_{X}||Q_{X})},$ (9)
$\eta_{\alpha}(P_{Y|X},\mathcal{P})=\sup_{P_{X}\in\mathcal{P}}\eta_{\alpha}(P_{X},P_{Y|X},\mathcal{P}).$ (10)
Here, $\mathcal{P}\subseteq\mathcal{P}_{\mathcal{X}}$ denotes a subset of the probability simplex on the alphabet $\mathcal{X}$. In this framework, the DPI states that $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})\leq 1$. We are then interested in the conditions for which $\eta_{\alpha}(P_{Y|X},\mathcal{P})<1$, that is, cases in which the inequality in (7) is strict. In this case, we say a strong data processing inequality (SDPI) holds. Similarly to the above, we define $\eta_{\text{TV}}(P_{Y|X},\mathcal{P})$ as the contraction coefficient of the total variation distance given that input distributions are picked from the set $\mathcal{P}$. In the most general case, an SDPI considers $\mathcal{P}=\mathcal{P}_{\mathcal{X}}$. However, we will later see that without any restriction, $\eta_{\alpha}(P_{Y|X})=1$ in many cases. We will therefore derive bounds for cases in which $\mathcal{P}\subsetneq\mathcal{P}_{\mathcal{X}}$. If the uncertainty set $\mathcal{P}$ is clear from context, we omit it for ease of notation.
III Conditions for $\eta_{\alpha}(P_{Y|X})=1$
@@
See Appendix A. ∎
@@
Here, $P$ is $3$-singular and $Q$ is $2$-singular. It is easy to see that the assertion in Proposition 3 applies to $Q$, but not to $P$.
IV Pinsker-type Inequalities for Rényi-divergence
@@
The following theorem bounds the $f_{\alpha}$-divergence as a function of the total variation between two distributions.
Theorem 6 (Reverse Pinkser’s Inequality for $f_{\alpha}$-divergence, special case of [1, Theorem 1]).
@@
$\Gamma_{\max}\coloneqq\sup_{P_{X},Q_{X}\in\mathcal{P},\,x\in\mathcal{X}}\frac{P_{X}(x)}{Q_{X}(x)},$ (16)
$\Gamma_{\min}\coloneqq\inf_{P_{X},Q_{X}\in\mathcal{P},\,x\in\mathcal{X}}\frac{P_{X}(x)}{Q_{X}(x)}.$ (17)
Then, the maximum $f_{\alpha}$-divergence corresponding to the $\alpha$-Rényi-divergence $D_{\alpha}(P_{X}||Q_{X})$ between the two distributions is upper bounded by
$D_{f_{\alpha}}(P_{X}||Q_{X})\leq\text{TV}(P_{X}||Q_{X})R_{\alpha}(\Gamma_{\max},\Gamma_{\min}),$ (18)
where
$R_{\alpha}(\Gamma_{\max},\Gamma_{\min})=\frac{\Gamma_{\max}^{\alpha}-1}{\Gamma_{\max}-1}-\frac{1-\Gamma_{\min}^{\alpha}}{1-\Gamma_{\min}}.$ (19)
We provide an alternative proof to [1]. Our proof utilizes the $E_{\gamma}$ identity of $f$-divergences.
@@
Fix $Q_{X}$ and $P_{X}$ and let $\Delta\coloneqq\text{TV}(P_{X}||Q_{X})$. The proof depends on the fact that $E_{\gamma}(P_{X}||Q_{X})$ can only be non-zero on the interval $(\Gamma_{\min},\Gamma_{\max})$. To see this, observe that
@@
This happens exactly when $\gamma\leq\min_{x}\frac{P_{X}(x)}{Q_{X}(x)}\eqqcolon\Gamma_{\min}(P_{X},Q_{X})$ or $\gamma\geq\max_{x}\frac{P_{X}(x)}{Q_{X}(x)}\coloneqq\Gamma_{\max}(P_{X},Q_{X})$. Hence, $E_{\gamma}$ is zero everywhere but on the interval $\gamma\in(\Gamma_{\min}(P_{X},Q_{X}),\Gamma_{\max}(P_{X},Q_{X}))\eqqcolon I_{\geq 0}$. From this, we get
$E_{\gamma}(P_{X}||Q_{X})=\begin{cases}0,&\text{ if }\gamma\notin I_{\geq 0},\\ u(\gamma),&\text{ if }\gamma\in I_{\geq 0}.\end{cases}$ (24)
What remains is to bound $u(\gamma)$ to obtain an upper bound in (6). To do this, note that $\gamma\mapsto E_{\gamma}(P||Q)$ is convex and increasing on $[0,1]$ and convex and decreasing on $(1,\infty)$ [10]. Hence, the maximum value is obtained at $\gamma=1$, where we have $E_{1}(Q_{X}||P_{X})=\Delta$. By convexity, we can then upper bound $u(\gamma)$ by the piecewise linear function
$u(\gamma)\leq g(\gamma)\coloneqq\begin{cases}0,&\text{if }\gamma\notin I_{\geq 0},\\ \Delta\frac{\gamma-\Gamma_{\min}(P_{X},Q_{X})}{1-\Gamma_{\min}(P_{X},Q_{X})},&\text{if }\gamma\in I_{\geq 0}\cap(0,1],\\ \Delta\frac{-\gamma+\Gamma_{\max}(P_{X},Q_{X})}{\Gamma_{\max}(P_{X},Q_{X})-1},&\text{if }\gamma\in I_{\geq 0}\cap(1,\infty).\end{cases}$ (25)
The theorem then follows from maximizing $g(\gamma)$ over $P_{X},Q_{X}\in\mathcal{P}$ and solving
@@
which yields a bound on the $f_{\alpha}$-divergence by (6). ∎
@@
The bound in Theorem 6 recovers the definition of Rényi-divergence of order infinity, that is, it is asymptotically tight with $D_{\infty}(P_{X}||Q_{X})=\log\Gamma_{\max}$. To see this, let again $\Delta\coloneqq\text{TV}(P_{X}||Q_{X})$ and observe that we can plug in the result form Theorem 6 into (4) to obtain
$D_{\alpha}(Q_{X}||P_{X})\leq\frac{1}{\alpha-1}\log\big{(}\Delta R_{\alpha}(\Gamma_{\max},\Gamma_{\min})+1\big{)}.$ (27)
Since $\Gamma_{\min}\leq 1$ and $\Gamma_{\max}\geq 1$, we get the limit as
$D_{\infty}(Q_{X}||P_{X})$ $\leq\lim_{\alpha\to\infty}\frac{1}{\alpha-1}\log\biggl{(}R_{\alpha}(\Gamma_{\max},\Gamma_{\min})\biggr{)}$ (28)
$=\lim_{\alpha\to\infty}\frac{\alpha}{\alpha-1}\log\Gamma_{\max}=\log\Gamma_{\max}.$ (29)
Further, the bound in [27, Theorem 35] can be obtained from Theorem 6 by assuming the lower bound $\Gamma_{\min}=0$.
Next, we present a generalized Pinsker’s inequality for $f_{\alpha}$-divergences of orders $\alpha>1$.
Theorem 8 (Pinsker’s inequality for Rényi-divergences).
@@
$g_{\alpha}(t)=\begin{cases}e^{2(\alpha-1)t^{2}}-1,&\text{ if }t<\nicefrac{{1}}{{\alpha}}\text{ and }\alpha<2,\\ (4t^{2}+1)^{\alpha-1}-1,&\text{ if }t<\nicefrac{{1}}{{\alpha}}\text{ and }\alpha\geq 2,\\ (1-t)^{1-\alpha}-1,&\text{ otherwise }\end{cases}$ (31)
@@
The proof follows from bounding the joint range [2] of the two divergences and is presented in Appendix B. ∎

### v3

# Bounds on the Privacy Amplification of Arbitrary Channels via the Contraction of $f_{\alpha}$-Divergence

@@ Abstract
We examine the privacy amplification of channels that do not necessarily satisfy any LDP guarantee by analyzing their contraction behavior in terms of $f_{\alpha}$-divergence, an $f$-divergence related to Rényi-divergence via a monotonic transformation. We present bounds on contraction for restricted sets of prior distributions via $f$-divergence inequalities and present an improved Pinsker’s inequality for $f_{\alpha}$-divergence based on the joint range technique by Harremoës and Vajda [1]. The presented bound is tight whenever the value of the total variation distance is larger than $\nicefrac{{1}}{{\alpha}}$. By applying these inequalities in a cross-channel setting, we arrive at strong data processing inequalities for $f_{\alpha}$-divergence that can be adapted to use-case specific restrictions of input distributions and channel. The application of these results to privacy amplification shows that even very sparse channels can lead to significant privacy amplification when used as a post-processing step after local differentially private mechanisms.
@@ I Introduction
An important property of local differential privacy (LDP) is its invariance to post-processing. Informally, the post-processing property states that whatever further processing is applied to the output of a system satisfying LDP will not weaken the overall system’s privacy guarantee. Mathematically, this property can be seen as a form of data processing inequality (DPI). As an extension of this observation, it is natural to ask about the privacy amplification of a post-processing operation. Basic results state that if a post-processing operation satisfying $\varepsilon_{2}$-LDP is applied to a system satisfying $\varepsilon_{1}$-LDP, the composition of both systems will satisfy at least $\min\{\varepsilon_{1},\varepsilon_{2}\}$-LDP. However, it is generally not know when arbitrary post-processing operations, that is, operations that themselves do not satisfy any LDP guarantee, may improve the overall systems LDP guarantee. We take steps to answer this question in this paper. In the context of data processing inequalities, this relates to the investigation of strong data processing inequalities (SDPIs), or the contraction of a channel with respect to a suitable measure of divergence (see [2, Chapter 33]). For pure LDP, the divergence measure of interest is the Rényi-divergence [3, 4] of order $\infty$, since
$\max_{y}\log\frac{P_{X|W=w}(y)}{P_{X|W=w^{\prime}}(y)}=\lim_{\alpha\to\infty}D_{\alpha}(P_{X|W=w}||P_{X|W=w^{\prime}}).$ (1)
Gilani et al. [5] suggest the notion of $\alpha$-Rényi-LDP as a relaxation of pure LDP, which bounds the maximum Rényi-divergence between the induced output distributions for some arbitrary order $\alpha$. With this framework at hand, studying the privacy amplification of a channel becomes equivalent to studying SDPIs for Rényi-divergences. While SDPIs have been extensively studied for $f$-divergences, results on Rényi-divergences are so far less explored. Although Rényi-divergence is not an $f$-divergence, it can be expressed as a monotonic transformation of a corresponding $f_{\alpha}$-divergence [6], a family of $f$-divergences parameterized by $\alpha$. This connection allows us to translate results on $f$-divergences to Rényi-divergences—results that are often easier to obtain.
I-A Contributions
In this work, we derive several technical tools for studying the privacy amplification of arbitrary discrete channels—those not necessarily satisfying any LDP guarantee—by analyzing how they affect the Rényi-LDP of composed systems. Our key contributions are:
  • We derive a general condition under which channels fail to contract Rényi-divergence, establishing structural limits on privacy amplification in Section III.
  • We develop new tools for bounding the $f_{\alpha}$-divergence in terms of total variation distance in Section IV. Specifically, we provide:
    • an improved and often tight version of Pinsker’s inequality for $f_{\alpha}$-divergence (for $\alpha>1$);
    • a simplification of the recently presented proof of Binette’s inequality [7] in [8] based on $E_{\gamma}$-divergence.
  • Using these tools, we analyze the contraction of $f_{\alpha}$-divergence and show how input distribution constraints can yield meaningful bounds in Section V.
  • Finally, in Section VI, we apply these bounds to the Rényi-LDP amplification of arbitrary post-processing channels. Our results show that many channels—despite not satisfying any (finite) LDP guarantee—can result in significant privacy amplification if applied to the output of a mechanism satisfying LDP. We illustrate these findings by applying the proposed bounds to randomized response mechanisms post-processed by channels with sparse channel matrices (i.e., matrices with many zeros).
I-B Related Work
@@
The contraction behavior of $f$-divergences has been extensively studied; see, e.g., [2, 9, 10, 11, 12, 13]. For Rényi-divergences, Polyanskiy and Wu [14] provide upper bounds in the Gaussian setting using the Hockeystick-divergence $E_{\gamma}$. Recently, Jin et al. [15] show that the standard total-variation-based upper bound fails for Rényi-divergence, though the universal $\chi^{2}$-based lower bound remains valid. In the context of privacy, Zamanlooy and Asoodeh [11] show that if a mechanism satisfies local differential privacy, strong data processing inequalities can be derived in terms of a channel’s LDP guarantee. These bounds are further applied to mixing times in [16]. Related, it is shown in [17] that the LDP guarantee of a mechanism is closely related to its $f$-divergence contraction properties. The $f_{\alpha}$-divergence framework enables the use of $f$-divergence inequalities in the Rényi setting, as developed in [6] and subsequent works, including an applications to SDPIs in [18].
@@ II Preliminaries
We use uppercase letters to denote random variables, lowercase letters to denote their realizations and caligraphic letters to denote sets. Specifically, we consider the two random variables $X$ and $Y$, where $X$ represents input data into a channel $P_{Y|X}$ (also referred to as mechanism if it satisfies an RLDP guarnatee in the following) that induces a random variable $Y$ at its output. Let $P_{XY}$ denote the joint distribution of $X$ and $Y$. Then we use $P_{XY}=P_{Y|X}\times P_{X}$ to imply that $P_{XY}(x,y)=P_{Y|X=x}(y)P_{X}(x)$ and $P_{Y}=P_{Y|X}\circ P_{X}$ to denote the marginalization $P_{Y}(y)=\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X}(x)$. We write $\mathcal{P}_{\mathcal{X}}$ for the probability simplex on the set $\mathcal{X}$. For $N\in\mathbb{N}$, we define $N]\coloneqq\{1,\dots,N\}$ as the set of positive integers up to $N$. Finally, $\log(\cdot)$ denotes the natural logarithm.
@@ II-A Rényi-divergence, $f_{\alpha}$-divergene and $E_{\gamma}$-divergence
We will use the Hockeystick-divergence $E_{\gamma}$ (from here on simply called $E_{\gamma}$-divergence) as a tool for obtaining bounds on the Rényi-divergence, as similarly done in [14, Section 3.2].
@@
can then be used to obtain any such $f$-divergence. For $\gamma=1$, the $E_{\gamma}$-divergence is equal to the total variation distance between the two distributions, i.e., $E_{1}(P||Q)=\text{TV}(P||Q)$.
@@ II-B Strong Data Processing Inequalities
All $f$-divergences (including the $f_{\alpha}$-divergence)—and hence also the Rényi-divergence—satisfy the data processing inequality (DPI). The DPI states that for any two probability measures $P_{X},Q_{X}$ on $\mathcal{X}$, and some transition kernel $P_{Y|X}$,
@@
We define the contraction coefficient $\eta_{\alpha}(P_{Y|X},\mathcal{P})$ of the $\alpha$-Rényi-divergence given a channel $P_{Y|X}$ as the maximum factor between the two sides of the above inequality, that is,
@@
$\coloneqq$ $\sup_{Q_{X}\in\mathcal{P}:D_{\alpha}(P_{X}||Q_{X})\neq 0}\frac{D_{\alpha}(P_{Y|X}\circ P_{X}||P_{Y|X}\circ Q_{X})}{D_{\alpha}(P_{X}||Q_{X})},$ (9)
$\eta_{\alpha}(P_{Y|X},\mathcal{P})\coloneqq\sup_{P_{X}\in\mathcal{P}}\eta_{\alpha}(P_{X},P_{Y|X},\mathcal{P}).$ (10)
Here, $\mathcal{P}\subseteq\mathcal{P}_{\mathcal{X}}$ denotes a subset of the probability simplex on the alphabet $\mathcal{X}$. In this framework, the DPI states that $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})\leq 1$. We are then interested in the conditions for which $\eta_{\alpha}(P_{Y|X},\mathcal{P})<1$, that is, cases in which the inequality in (7) is strict. In this case, we say a strong data processing inequality (SDPI) holds. Similarly to the above, we define $\eta_{\text{TV}}(P_{Y|X},\mathcal{P})$ as the contraction coefficient of the total variation distance given that input distributions are picked from the set $\mathcal{P}$. In the most general case, an SDPI considers $\mathcal{P}=\mathcal{P}_{\mathcal{X}}$. However, we will see below that without any restriction, $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$ in many cases. We will therefore derive bounds for cases in which $\mathcal{P}\subsetneq\mathcal{P}_{\mathcal{X}}$.
III Conditions for $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$
@@
We provide a constructive proof in Appendix B. Here, we show Proposition 3 by a simple chain of arguments following from previous results. Let $(X,Y)\sim P_{Y|X}\times P_{X}$ be arbitrary but fixed. Define the maximal correlation as
$\rho_{m}(X;Y)\coloneqq\sup_{f,g}\mathbb{E}_{(X,Y)\sim P_{XY}}\Big{[}f(X)g(Y)\Big{]},$ (12)
where the supremum is over all $f:\mathcal{X}\to\mathbb{R}$ and $g:\mathcal{Y}\to\mathbb{R}$ such that
$\mathbb{E}\big{[}f(X)\big{]}=\mathbb{E}\big{[}g(Y)\big{]}=0,\quad\mathbb{E}\big{[}f^{2}(X)\big{]}=\mathbb{E}\big{[}g^{2}(Y)\big{]}=1.$ (13)
The contraction coefficient of the $\chi^{2}$-divergence is closely related to $\rho_{m}$, in particular, it is shown in, e.g., [2, Theorem 33.6(c)] that
$\eta_{\chi^{2}}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=\sup_{P_{X}\in\mathcal{P}_{\mathcal{X}}}\rho_{m}^{2}(X;Y)\eqqcolon s(P_{Y|X}).$ (14)
It was shown by Ahlswede and Gács [13] that $s(P_{Y|X})<1$ if and only if for every pair $x,x^{\prime}$, it holds that $\textnormal{supp}(P_{Y|X=x})\cap\textnormal{supp}(P_{Y|X=x^{\prime}})\neq\emptyset$. Together with the data processing inequality for $f$-divergences, this implies that whenever (11) holds, we have $\eta_{\chi^{2}}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$. Next, it was shown in [15, Corollary 2] that
$\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})\geq\eta_{\chi^{2}}(P_{Y|X},\mathcal{P}_{\mathcal{X}}).$ (15)
By the data processing inequality for Rényi-divergences, we also have $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})\leq 1$. Hence, if (11) holds, we have $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$. ∎
@@
Here, $P$ is 33-singular and $Q$ is 22-singular. It is easy to see that the assertion in Proposition 3 applies to $Q$, but not to $P$.
IV Pinsker-type Inequalities for $f_{\alpha}$-divergence
@@
The following theorem bounds the $f_{\alpha}$-divergence as a function of the total variation between two distributions. It was first shown by Binette [7] for general $f$-divergences. Here, we present a proof that utilizes the integral representation of $f$-divergences via the $E_{\gamma}$-divergence in (6). We remark that the same technique was recently used to proof Binette’s inequality in a quantum context by Hirche and Tomamichel [8, Proposition 5.2]. However, the proof presented below slightly simplifies arguments, and we include it here as we believe it to be instructive.
Theorem 6 (Reverse Pinkser’s Inequality for $f_{\alpha}$-divergence, special case of [7, Theorem 1]).
@@
$\Gamma_{\max}(X)\coloneqq\sup_{P_{X},Q_{X}\in\mathcal{P},\,x\in\mathcal{X}}\frac{P_{X}(x)}{Q_{X}(x)},$ (20)
$\Gamma_{\min}(X)\coloneqq\inf_{P_{X},Q_{X}\in\mathcal{P},\,x\in\mathcal{X}}\frac{P_{X}(x)}{Q_{X}(x)}.$ (21)
Then, the maximum $f_{\alpha}$-divergence between the two distributions is upper bounded by
$D_{f_{\alpha}}(P_{X}||Q_{X})\leq\textnormal{TV}(P_{X}||Q_{X})R_{\alpha}(\Gamma_{\max}(X),\Gamma_{\min}(X)),$ (22)
where for $u\in[1,\infty)\cup\{\infty\}$ and $v\in[0,1]$,
$R_{\alpha}(u,v)=\frac{u^{\alpha}-1}{u-1}-\frac{1-v^{\alpha}}{1-v}.$ (23)
@@
Fix $Q_{X}$ and $P_{X}$ and let $\Delta\coloneqq\text{TV}(P_{X}||Q_{X})$. The proof depends on the fact that the only interval on which we may have $E_{\gamma}(P_{X}||Q_{X})\neq 0$ is $(\Gamma_{\min}(X),\Gamma_{\max}(X))$. To see this, observe that
$E_{\gamma}(P_{X}||Q_{X})=\frac{1}{2}\sum_{x\in\mathcal{X}}\biggl{|}P_{X}(x)-\gamma Q_{X}(x)\biggr{|}-\frac{1}{2}|1-\gamma|,$ (24)
@@
This happens exactly when
$\gamma\leq\min_{x}\frac{P_{X}(x)}{Q_{X}(x)}\eqqcolon\Gamma_{\min}(X,P_{X},Q_{X}),$ (28)
or,
$\gamma\geq\max_{x}\frac{P_{X}(x)}{Q_{X}(x)}\eqqcolon\Gamma_{\max}(X,P_{X},Q_{X}).$ (29)
Hence, $E_{\gamma}$ must be zero everywhere but on the interval between these two quantities, that is, the only values of $\gamma$ that allow a non-zero value of $E_{\gamma}$ are
$\gamma\in(\Gamma_{\min}(X,P_{X},Q_{X}),\Gamma_{\max}(X,P_{X},Q_{X}))\eqqcolon I_{\geq 0}.$ (30)
From this, we get
@@
What remains is to bound $E_{\gamma}$ on the interval $I_{\geq 0}$, which yields an upper bound on $D_{f_{\alpha}}$ by (6). To do this, note that $\gamma\mapsto E_{\gamma}(P||Q)$ is convex and increasing on $[0,1]$ and convex and decreasing on $(1,\infty)$ [14]. Hence, the maximum value is obtained at $\gamma=1$, where we have $E_{1}(Q_{X}||P_{X})=\Delta$. By convexity, we can then upper bound $E_{\gamma}$ by the piecewise linear function
$E_{\gamma}$ $(P_{X}||Q_{X})\leq g(\gamma)$ (32)
$\coloneqq\begin{cases}0,&\text{if }\gamma\notin I_{\geq 0},\\[6.99997pt] \Delta\frac{\gamma-\Gamma_{\min}(X,P_{X},Q_{X})}{1-\Gamma_{\min}(X,P_{X},Q_{X})},&\text{if }\gamma\in I_{\geq 0}\cap(0,1],\\[6.99997pt] \Delta\frac{-\gamma+\Gamma_{\max}(X,P_{X},Q_{X})}{\Gamma_{\max}(X,P_{X},Q_{X})-1},&\text{if }\gamma\in I_{\geq 0}\cap(1,\infty).\end{cases}$ (33)
The theorem then follows from maximizing $g(\gamma)$ over $P_{X},Q_{X}\in\mathcal{P}$, which together with (6) yields
$D_{f_{\alpha}}(P_{X}||Q_{X})\leq\alpha(\alpha-1)\int_{\Gamma_{\min}(X)}^{\Gamma_{\max}(X)}g(\gamma)\gamma^{\alpha-2}d\gamma,$ (34)
and the desired statement follows by solving this integral. ∎
@@
The bound in Theorem 6 recovers the definition of Rényi-divergence of order infinity, that is, it is asymptotically tight with $D_{\infty}(P_{X}||Q_{X})=\log\Gamma_{\max}(X)$. To see this, let again $\Delta\coloneqq\text{TV}(P_{X}||Q_{X})$ and observe that we can plug in the result form Theorem 6 into (4) to obtain
$D_{\alpha}(Q_{X}||P_{X})\leq\frac{1}{\alpha-1}\log\big{(}\Delta R_{\alpha}(\Gamma_{\max}(X),\Gamma_{\min}(X))+1\big{)}.$ (35)
Since $\Gamma_{\min}(X)\leq 1$ and $\Gamma_{\max}(X)\geq 1$, we get the limit as
$D_{\infty}(Q_{X}||P_{X})$ $\leq\lim_{\alpha\to\infty}\frac{1}{\alpha-1}\log\biggl{(}R_{\alpha}\big{(}\Gamma_{\max}(X),\Gamma_{\min}(X)\big{)}\bigg{)}$ (36)
$=\lim_{\alpha\to\infty}\frac{\alpha}{\alpha-1}\log\Gamma_{\max}(X)=\log\Gamma_{\max}(X).$ (37)
The bound in [6, Theorem 35], [25, Theorem 3] is obtained from Theorem 6 by assuming the lower bound $\Gamma_{\min}(X)=0$.
Next, we present a generalized Pinsker’s inequality for $f_{\alpha}$-divergence of order $\alpha>1$.
Theorem 8 (Pinsker’s inequality for $f_{\alpha}$-divergences).
@@
[…]
@@
The proof follows from bounding the joint range [1] of the divergences and is presented in Appendix A. ∎